OpenStudy (anonymous):

Use the properties of exponents to solve the equation (51/3)x = 5.

OpenStudy (diamondboy):

multiply both sides by 3

OpenStudy (thomas5267):

Is the question \(\left(\frac{51}{3}\right)^x=5\)?

OpenStudy (arabpride):

The equation does not make sense....